I need your help
to solve for x
in the equation
3ln(2-x)=6
I tried to solve it but I'm not sure if my answer is correct
3ln(2-x)=6
ln2/lnx=2
ln2=2lnx
e^(2-x)=2
ln(2-x)=2
e^2=e^ln(x+5)
e^2-5=x

Answer by venugopalramana(3286)

I need your help
to solve for x
in the equation
3ln(2-x)=6.......OK
I tried to solve it but I'm not sure if my answer is correct
3ln(2-x)=6............OK
ln2/lnx=2....NO..NO...NO..THERE IS NO SUCH FORMULA..LN(X/Y)=LN(X)-LN(Y)..
BUT LN(X-Y) IS NOT EQUAL TO LN(X-Y)
SO LN(2-X)=6/3=2
E^2=2-X
X=2-E^2